OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 18 and a15 = -38 ?

OpenStudy (helder_edwin):

is the sequence is arithmetic then the terms (from the second) would be \[\large a_2=a_1+d \] \[\large a_3=a_2+d=a_1+2d \] \[\large a_4=a_3+d=a_1+3d \] so in general the n-th term would be \[\large a_n=a_{n-1}+d=a_1+(n-1)d \]

OpenStudy (helder_edwin):

u have to compute the value of \(d\)

OpenStudy (anonymous):

howw?

OpenStudy (anonymous):

@helder_edwin

OpenStudy (helder_edwin):

u have a1 and a15, so \[\large a_{15}=a_{14}+d=a_1+14d \]

OpenStudy (helder_edwin):

solve for d

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

-138?

OpenStudy (helder_edwin):

what is that?

OpenStudy (anonymous):

umm the answer?

OpenStudy (helder_edwin):

u mean d=-138 or a41=-138

OpenStudy (anonymous):

a41

OpenStudy (helder_edwin):

let's see \[\large a_{15}=-38\qquad a_1=18 \] so \[\large -38=18+14d\Rightarrow d=-4 \] so \[\large a_{41}=a_1+40d=18+40(-4)=-142 \]

OpenStudy (anonymous):

damm

OpenStudy (anonymous):

thanks

OpenStudy (helder_edwin):

u r welcome